# Feature Intensity Mapping: Polycyclic Aromatic Hydrocarbon Emission from All Galaxies Across Cosmic Time

### v1

@@ 4.1 Feature Models
The intensity mapping analysis employs features from the 1C starburst template presented by T. S. Y. Lai et al. (2020), which is derived from combined AKARI and Spitzer observations, ranging from 5 to 38 m. We obtained the relative luminosities of these features, including both emission lines and PAH bands, derived from the spectral decomposition of the 1C template at a resolution of 100. Since this spectral resolution may not fully resolve these emission lines and bands, we implemented several assumptions to generate more realistic intrinsic profiles in our model: (i) delta functions for narrow emission lines, and (ii) width corrections for broad dust features (e.g., PAH features) based on JWST’s high spectral resolution spectroscopy of the starburst ring in NGC 7469 (T. S. Y. Lai et al., 2022, 2023). The width correction shrinks the Drude profiles’ FWHM by a factor of 2–4 for the shorter wavelength PAH features (6.22 m and below), while the longer wavelength profiles are already well-resolved by Spitzer. Using the updated FWHM values for each PAH feature, we derived the intrinsic PAH profiles by conserving flux under a Drude profile, following the approach of J. D. T. Smith et al. (2007). We further group blended PAH features into various complexes at 7.7, 11.3, 12.7, and 17 m, with the individual components comprising each complex listed in Table 3 of J. D. T. Smith et al. (2007).
@@ 4.2 Luminosity Density and Bias
and the SFR–halo mass () relation from P. S. Behroozi et al. (2013a, b), who derived an empirical model for the average galaxy SFR as a function of and , , calibrated to observations. We use their publicly released888sfh_z0_z8.tar.gz file from https://www.peterbehroozi.com/data.html model of over a range of and bins, and interpolate between these bins to construct a continuous model. With these ingredients, along with the feature luminosity ratios listed in Table 1, we obtain the mean feature luminosity as a function of halo mass and redshift, . We then estimate the luminosity density and large-scale bias by integrating over the halo mass:
@@
In reality, when the observing channels are able to resolve a spectral feature across multiple channels, the observed intensity in a given channel will be the integrated contribution from a range of redshifts, weighted by the filter profile. This is effectively equivalent to applying a kernel (with the shape determined by the filter profile) to the bias-weighted intensity shown in the lower panels of Figures 4 and 5. Since this quantity vary smoothly with redshift, we expect that the actual observed bias-weighted intensity within each spectral channel will be well approximated by our simplified calculation. While the per pixel noise rms (Figure 1 and 2) is orders of magnitude higher than the expected signal in individual pixels (lower panel of Figures 4 and 5), our constraints on the aggregate extragalactic emission are derived through 3D clustering. As will be demonstrated in Section 5, the large-scale correlations significantly suppress the noise, allowing for strong constraints on these emission features.
@@ 4.3 Power Spectrum
Using the spectral feature model described in Sections 4.1 and 4.2, along with the observing filters defined in Section 3, we compute the cross power spectrum matrices for SPHEREx, PHI, and FIRESS using Equation (7).
Figures 6 and 7 show the clustering power spectrum matrix at the lowest- bin for SPHEREx and PRIMA, respectively. The diagonal elements represent the auto-power spectra of individual spectral channels. The band along the diagonal reflects the cross-correlations between nearby channels, which are expected to be strong due to the broad spectral width of features that span multiple channels. Additional off-diagonal streaks correspond to correlations between widely separated features that trace the same redshift, leading to coherence between non-adjacent channels.
@@ 5.1 SPHEREx
We find that, despite the very different depth and survey area between the all-sky and deep-field configurations, the on detecting for the features is of a similar order of magnitude, with slightly higher sensitivity expected in the deep fields. In practice, image systematics and foregrounds are better studied in the deep fields than in the all-sky maps for SPHEREx, making the deep fields preferable for conducting this measurement compared to the all-sky survey.
@@ 5.2 PRIMA
With 1,000 hours of integration time for both PHI and FIRESS, PHI achieves values approximately twice as high as FIRESS for all PAH features. The 7.7 and 11.3 m PAHs have the highest , as they are relatively brighter PAH features in our SED model. The next highest total sensitivities come from the 12.7 and 17 m features. Although these are not as intrinsically bright as, for example, the 6.22 m feature, they are observed at lower redshifts, which boosts their observed intensities. Our forecast predicts that the aggregate emission from multiple PAH features across a broad redshift range, spanning from to , can be detected with high significance through their large-scale clustering using our feature intensity mapping technique.
@@ 6.1 Intensity Mapping with Unresolved Sources
In our model described in Section 4, we consider the emission from all sources within the survey. However, brighter sources are expected to be individually detectable in these surveys. If the detection threshold is sufficiently deep such that most of the total emission is captured by resolved sources, performing intensity mapping becomes less meaningful.
@@
A recent study by E. Zhang et al. (2025) investigated the individual detectability of the 3.3 m PAH feature in both the SPHEREx all-sky and deep-field configurations using the SPHEREx Sky Simulator (B. P. Crill et al., 2025). They derived the SFR and -band magnitude limits above which the 3.3 m PAH feature can be individually detected. Assuming that such detected sources can be masked in the data, we estimate the bias-weighted intensity contributed only by sources below the detection threshold, and repeat our sensitivity forecasts to evaluate the detectability of the remaining unresolved emission. Specifically, we adopt the SFR limits on SPHEREx all-sky configuration from E. Zhang et al. (2025), and assume the deep fields SFR limit to be eight times lower, corresponding to the approximated point source sensitivity improvement on deep fields compared to all sky. In this estimate, we neglect the potential signal loss and mode coupling effects introduced by masking.
@@
The bias-weighted intensity from only the sources below the detection threshold is shown in Figure 12. We find that, for SPHEREx, at and in the all-sky and deep-field configurations, respectively, a significant fraction of the 3.3 m PAH emission originates from galaxies that can be individually detected. Beyond these redshift limits, however, the majority of the emission comes from faint, unresolved sources below the detection limit. Figure 13 displays the on detecting the signals from only the sub-threshold sources. Our method can still achieve a detection of the aggregate 3.3 m PAH intensity at these redshifts. This highlights the complementarity of our approach to individual source detection techniques.
@@
For the case of PRIMA, there is no study analogous to E. Zhang et al. (2025) that quantifies the PAH detection threshold for individual galaxies. Using a semi-analytic model and considering 1,500 hours of PHI observations, L. Bisigello et al. (2024) estimated the IR luminosity and SFR distributions of detectable galaxies. Their forecast is performed on both a deep (1 deg2) and a wide (10 deg2) survey, each with the same 1,500 hours integration time. Our fiducial PHI case of 1 deg2 with 1,000 hour integration is similar to their deep survey with a slightly shorter integration, and thus their detection limit can serve as a lower bound of the detectable IR luminosity and the SFR values. Applying their SFR limits to our model of PAH intensity, we find that at , the majority of the PAH total intensity is contributed by faint, unresolved galaxies. As shown in Section 5.2, applying our feature intensity mapping technique to PHI can achieve for multiple PAH features at this redshift range. These aggregate signals from faint sources will be uniquely accessible through intensity mapping, offering a complementary capability to individual galaxy detection with PRIMA.
@@ 6.2 Survey Optimization
As PRIMA is still in the design phase and a concrete survey plan has not yet been finalized, we explore the sensitivity of PAH intensity mapping performance under different survey configurations. Specifically, assuming a fixed total integration time of 1,000 hours as our fiducial case, we vary the survey area, which effectively changes the noise level and the effective number of modes (through the factor in Equation (10)). We repeat our sensitivity forecasts for scenarios where the survey area is scaled up and down by a factor of ten relative to our fiducial PHI and FIRESS setups. Since most of the sensitivity comes from scales much smaller than the survey size, we fix the range and binning when scaling the survey area, and scale the number of modes (Equation 10) proportionally to the survey area. The resulting total sensitivity for each PAH feature under these configurations is summarized in Table 2. We find that our adopted fiducial survey sizes are close to optimal for both PHI and FIRESS.
@@
10.3 | 14.4 | 9.19 | 5.06 | 6.34 | 4.03 | |
1.41 | 2.64 | 1.46 | 0.18 | 0.57 | 0.16 | |
40.3 | 61.5 | 40.1 | 21.1 | 31.1 | 20.9 | |
6.30 | 8.62 | 4.64 | 2.86 | 3.25 | 1.39 | |
13.9 | 17.6 | 10.1 | 7.03 | 6.98 | 4.04 | |
65.8 | 68.2 | 39.3 | 35.7 | 31.1 | 15.9 | |
13.6 | 14.7 | 7.86 | 5.29 | 5.62 | 2.84 | |
35.3 | 35.8 | 20.4 | 14.7 | 13.9 | 8.03 | |
24.6 | 28.2 | 19.9 | 11.1 | 10.0 | 5.89 |
Note. — The for the nine PAH features for PHI and FIRESS with 1,000 hours of integration time under different survey area assumptions. The middle case represents our fiducial survey setup, while the other two cases correspond to survey areas that are scaled by a factor of 10 smaller and larger, respectively.
@@ 6.3 Implication of PAH Intensity Mapping
Using the feature intensity mapping technique developed in this work, we are able to perform 3D intensity mapping for broad spectral features such as PAHs, analogous to line intensity mapping. Our forecasts demonstrate promising prospects for probing multiple PAH features across cosmic time. SPHEREx, which covers wavelengths up to 5 m, will enable mapping of the 3.3 m PAH feature up to , while PRIMA is expected to be capable of mapping multiple PAH features with high significance up to .
@@
In addition to auto- and cross-correlations within the same survey, synergy with external datasets provides powerful avenues for probing galaxy physics in the cosmic background. For example, by cross-correlating SPHEREx channel maps with galaxy catalogs, one can extract the aggregate galaxy SEDs, including continuum emission and spectral features such as lines and PAHs, as a function of wavelength and redshift (Y.-T. Cheng & T.-C. Chang, 2022). Furthermore, joint analyses of the PAH background mapped by SPHEREx or PRIMA with LIM surveys targeting different spectral lines (e.g., [C II] or CO) can give a more comprehensive census of galaxy physics by simultaneously probing multiple phases of the ISM through distinct tracers (G. Sun et al., 2019).
In addition, similar to line intensity mapping, PAHs can serve as tracers of the 3D density field, providing a novel probe of large-scale structure at redshifts well beyond the reach of traditional galaxy surveys.
@@ 6.4 Future Work
The feature intensity mapping technique developed in this work is timely for the recently launched SPHEREx mission and the proposed PRIMA mission, both of which will conduct large-scale spectral imaging surveys in the infrared, where multiple PAH features fall within their wavelength coverage.
@@
This work is the third in our series on performing inference using the full covariance of spectral imaging surveys, building upon Y.-T. Cheng et al. (2023) and Y.-T. Cheng et al. (2024). In Y.-T. Cheng et al. (2023), we focused on reconstructing the principal modes of the continuum from cross-correlations in broadband photometric intensity maps. In Y.-T. Cheng et al. (2024), we explored the application of the same method for line intensity mapping. In this work, we extend our model to include broad spectral features such as PAHs.
@@ 7 Conclusion
For PRIMA, we consider nine PAH features spanning to m and redshifts . We assume PHI and FIRESS surveys covering 1 and 0.1 deg2, respectively, each with 1,000 hours of integration. Our forecasts show that multiple PAH features can achieve detections of their bias-weighted intensity per bins of redshifts at , with the two bright PAH features (7.7 and 11.3 m) detectable out to .
While individually detected galaxies at high redshift will be biased toward the brightest sources, our results demonstrate that IM offers a powerful complementary approach to recover the integrated emission from the full galaxy population, including faint and unresolved sources. Our feature intensity mapping technique enables the extension of 3D intensity mapping, previously limited to narrow emission lines with LIM, to general broad spectral features such as PAHs. This is especially crucial given the importance of PAHs in tracing star formation, dust and metallicity evolution in galaxies. The development of this technique is timely: SPHEREx has just begun its survey, and, if selected for implementation, PRIMA could launch early in the next decade.

### v2

@@ 4.1 Feature Models
The IM analysis employs features from the 1C starburst template presented by T. S. Y. Lai et al. (2020), which is derived from combined AKARI and Spitzer observations, ranging from 5 to 38 m. We obtained the relative luminosities of these features, including both emission lines and PAH bands, derived from the spectral decomposition of the 1C template at a resolution of 100. Since this spectral resolution may not fully resolve these emission lines and bands, we implemented several assumptions to generate more realistic intrinsic profiles in our model: (i) delta functions for narrow emission lines, and (ii) width corrections for broad dust features (e.g., PAH features) based on JWST’s high spectral resolution spectroscopy of the starburst ring in NGC 7469 (T. S. Y. Lai et al., 2022, 2023). The width correction shrinks the Drude profiles’ FWHM by a factor of 2–4 for the shorter wavelength PAH features (6.22 m and below), while the longer wavelength profiles are already well-resolved by Spitzer. Using the updated FWHM values for each PAH feature, we derived the intrinsic PAH profiles by conserving flux under a Drude profile, following the approach of J. D. T. Smith et al. (2007). We further group blended PAH features into various complexes at 7.7, 11.3, 12.7, and 17 m, with the individual components comprising each complex listed in Table 3 of J. D. T. Smith et al. (2007).
@@
For the SPHEREx case, in addition to PAHs, we also consider two other spectral features at 3.4 m and 3.48 m, as well as the nearby 4.06 m Br line. While PAH features arise from aromatic C–H stretching modes, the 3.4 m feature originates from aliphatic C–H stretches, typically associated with carbonaceous chain structures. The broad 3.48 m plateau feature is often attributed to aromatic carriers because it shows a slightly tighter correlation with the 3.3 m PAH band than with the 3.4 m aliphatic band (P. Pilleri et al., 2015; T. S. Y. Lai et al., 2020). However, this difference is only marginal, and further studies are needed to constrain the true origin of the 3.48 m plateau feature.
@@ 4.2 Luminosity Density and Bias
which is derived from a sample of star-forming galaxies. The PAH intensity is expected to be dominated by star-forming population, while the contribution from quiescent galaxies will be negligible. This is because the excitation of PAH molecules requires the absorption of far-ultraviolet photons with energies between and eV, typically originating from regions of active star formation (A. Leger & J. L. Puget, 1984). Observational studies with Spitzer have demonstrated that passively evolving early-type galaxies exhibiting detectable PAH features are rare, constituting of the population (J. N. Bregman et al., 2006; A. Bressan et al., 2006). Therefore, their contribution to the total PAH emission is expected to be negligible.
We adopt the SFR–halo mass () relation from P. S. Behroozi et al. (2013a, b), who derived an empirical model for the average galaxy SFR as a function of and , , calibrated to observations. We use their publicly released888sfh_z0_z8.tar.gz file from https://www.peterbehroozi.com/data.html model of over a range of and bins, and interpolate between these bins to construct a continuous model. With these ingredients, along with the feature luminosity ratios listed in Table 1, we obtain the mean feature luminosity as a function of halo mass and redshift, . We then estimate the luminosity density and large-scale bias by integrating over the halo mass:
@@
In reality, when the observing channels are able to resolve a spectral feature across multiple channels, the observed intensity in a given channel will be the integrated contribution from a range of redshifts, weighted by the filter profile. This is effectively equivalent to applying a kernel (with the shape determined by the filter profile) to the bias-weighted intensity shown in the lower panels of Figures 4 and 5. Since this quantity varies smoothly with redshift, we expect that the actual observed bias-weighted intensity within each spectral channel will be well approximated by our simplified calculation. While the per pixel noise rms (Figure 1 and 2) is orders of magnitude higher than the expected signal in individual pixels (lower panel of Figures 4 and 5), our constraints on the aggregate extragalactic emission are derived through 3D clustering. As will be demonstrated in Section 5, the large-scale correlations significantly suppress the noise, allowing for strong constraints on these emission features.
@@ 4.3 Power Spectrum
Using the spectral feature model described in Sections 4.1 and 4.2, along with the observing filters defined in Section 3, we compute the cross-power spectrum matrices for SPHEREx, PHI, and FIRESS using Equation (7).
Figures 6 and 7 show the clustering power spectrum matrix at the lowest- bin for SPHEREx and PRIMA, respectively. The diagonal elements represent the auto-power spectra of individual spectral channels. The band along the diagonal reflects the cross correlations between nearby channels, which are expected to be strong due to the broad spectral width of features that span multiple channels. Additional off-diagonal streaks correspond to correlations between widely separated features that trace the same redshift, leading to coherence between non-adjacent channels.
@@ 5.1 SPHEREx
We find that, despite the very different depths and survey areas between the all-sky and deep-field configurations, the on detecting for the features is of a similar order of magnitude, with slightly higher sensitivity expected in the deep fields. In practice, image systematics and foregrounds are better studied in the deep fields than in the all-sky maps for SPHEREx, making the deep fields preferable for conducting this measurement compared to the all-sky survey.
@@ 5.2 PRIMA
With 1,000-hr of integration time for both PHI and FIRESS, PHI achieves values approximately twice as high as FIRESS for all PAH features. The 7.7 and 11.3 m PAHs have the highest because they are relatively brighter PAH features in our SED model. The next highest total sensitivities come from the 12.7 and 17 m features. Although these are not as intrinsically bright as, for example, the 6.22 m feature, they are observed at lower redshifts, which boosts their observed intensities. Our forecast predicts that the aggregate emission from multiple PAH features across a broad redshift range, spanning from to , can be detected with high significance through their large-scale clustering using our feature intensity mapping technique.
@@ 6.1 Intensity Mapping with Unresolved Sources
In our model described in Section 4, we consider the emission from all sources within the survey. However, brighter sources are expected to be individually detectable in these surveys. If the detection threshold is sufficiently deep such that most of the total emission is captured by resolved sources, performing IM becomes less meaningful.
@@
A recent study by E. Zhang et al. (2025) investigated the individual detectability of the 3.3 m PAH feature in both the SPHEREx all-sky and deep-field configurations using the SPHEREx Sky Simulator (B. P. Crill et al., 2025). They derived the SFR and -band magnitude limits above which the 3.3 m PAH feature can be individually detected. Assuming that such detected sources can be masked in the data, we estimate the bias-weighted intensity contributed only by sources below the detection threshold, and repeat our sensitivity forecasts to evaluate the detectability of the remaining unresolved emission. Specifically, we adopt the SFR limits on SPHEREx all-sky configuration from E. Zhang et al. (2025), and assume the deep-fields SFR limit to be eight times lower, corresponding to the approximated point source sensitivity improvement on deep fields compared to all sky. In this estimate, we neglect the potential signal loss and mode coupling effects introduced by masking.
@@
The bias-weighted intensity from only the sources below the detection threshold is shown in Figure 12. We find that, for SPHEREx, at and in the all-sky and deep-field configurations, respectively, a significant fraction of the 3.3 m PAH emission originates from galaxies that can be individually detected. Beyond these redshift limits, however, the majority of the emission comes from faint, unresolved sources below the detection limit. Figure 13 displays the on detecting the signals from only the subthreshold sources. Our method can still achieve a detection of the aggregate 3.3 m PAH intensity at these redshifts. This highlights the complementarity of our approach to individual source detection techniques.
@@
For the case of PRIMA, there is no study analogous to E. Zhang et al. (2025) that quantifies the PAH detection threshold for individual galaxies. Using a semi-analytic model and considering 1,500 hr of PHI observations, L. Bisigello et al. (2024) estimated the IR luminosity and SFR distributions of detectable galaxies. Their forecast is performed on both a deep (1 deg2) and a wide (10 deg2) survey, each with the same 1,500 hr integration time. Our fiducial PHI case of 1 deg2 with 1,000 hr integration is similar to their deep survey with a slightly shorter integration, and thus their detection limit can serve as a lower bound of the detectable IR luminosity and the SFR values. Applying their SFR limits to our model of PAH intensity, we find that at , the majority () of the PAH total intensity is contributed by faint, unresolved galaxies. As shown in Section 5.2, applying our feature intensity mapping technique to PHI can achieve for multiple PAH features at this redshift range. These aggregate signals from faint sources will be uniquely accessible through IM, offering a complementary capability to individual galaxy detection with PRIMA.
@@ 6.2 Survey Optimization
Since PRIMA is still in the design phase and a concrete survey plan has not yet been finalized, we explore the sensitivity of PAH intensity mapping performance under different survey configurations. Specifically, assuming a fixed total integration time of 1,000 hr as our fiducial case, we vary the survey area, which effectively changes the noise level and the effective number of modes (through the factor in Equation (10)). We repeat our sensitivity forecasts for scenarios where the survey area is scaled up and down by a factor of 10 relative to our fiducial PHI and FIRESS setups. Since most of the sensitivity comes from scales much smaller than the survey size, we fix the range and binning when scaling the survey area, and scale the number of modes (Equation 10) proportionally to the survey area. The resulting total sensitivity for each PAH feature under these configurations is summarized in Table 2. We find that our adopted fiducial survey sizes are close to optimal for both PHI and FIRESS.
@@
| 10.3 | 14.4 | 9.19 | 5.06 | 6.34 | 4.03 | |
| 1.41 | 2.64 | 1.46 | 0.18 | 0.57 | 0.16 | |
| 40.3 | 61.5 | 40.1 | 21.1 | 31.1 | 20.9 | |
| 6.30 | 8.62 | 4.64 | 2.86 | 3.25 | 1.39 | |
| 13.9 | 17.6 | 10.1 | 7.03 | 6.98 | 4.04 | |
| 65.8 | 68.2 | 39.3 | 35.7 | 31.1 | 15.9 | |
| 13.6 | 14.7 | 7.86 | 5.29 | 5.62 | 2.84 | |
| 35.3 | 35.8 | 20.4 | 14.7 | 13.9 | 8.03 | |
| 24.6 | 28.2 | 19.9 | 11.1 | 10.0 | 5.89 |
Note. — The for the nine PAH features for PHI and FIRESS with 1,000 hr of integration time under different survey area assumptions. The middle case represents our fiducial survey setup, while the other two cases correspond to survey areas that are scaled by a factor of 10 smaller and larger, respectively.
@@ 6.3 Implication of PAH Intensity Mapping
Using the feature intensity mapping technique developed in this work, we are able to perform 3D intensity mapping for broad spectral features such as PAHs, analogous to LIM. Our forecasts demonstrate promising prospects for probing multiple PAH features across cosmic time. SPHEREx, which covers wavelengths up to 5 m, will enable mapping of the 3.3 m PAH feature up to , while PRIMA is expected to be capable of mapping multiple PAH features with high significance up to .
@@
Although the detectability of the 3.4 m aliphatic feature with SPHEREx is marginal with our model, a constraint on the aliphatic-to-aromatic ratio (traced by the 3.4 and 3.3 m intensity ratio) across a broad redshift range would enable studies of the evolution of the ISM. In particular, this ratio constrains the relative abundances of aromatic and aliphatic bonds, which are key indicators of the carbon budget and its evolution over cosmic time. The aliphatic-to-aromatic ratio is observed to decrease with increasing intensity of the far-ultraviolet radiation field (C. Joblin et al., 1996; P. Pilleri et al., 2015), since aliphatic bonds are more susceptible to destruction through photon-driven processing. Previous studies of the aliphatic-to-aromatic ratio outside the Milky Way have been largely limited to nearby galaxies (e.g., M. Yamagishi et al., 2012; T. S. Y. Lai et al., 2020). With the advent of JWST, however, this diagnostic can now be extended to higher redshifts (; J. Lyu et al., 2025). Our feature intensity mapping technique provides a complementary probe to the averaged measurements from all galaxies, while JWST focuses on individual targets.
In addition to auto and cross correlations within the same survey, synergy with external datasets provides powerful avenues for probing galaxy physics in the cosmic background. For example, by cross-correlating SPHEREx channel maps with galaxy catalogs, one can extract the aggregate galaxy SEDs, including continuum emission and spectral features such as lines and PAHs, as a function of wavelength and redshift (Y.-T. Cheng & T.-C. Chang, 2022). Furthermore, joint analyses of the PAH background mapped by SPHEREx or PRIMA with LIM surveys targeting different spectral lines (e.g., [C II] or CO) can give a more comprehensive census of galaxy physics by simultaneously probing multiple phases of the ISM through distinct tracers (G. Sun et al., 2019).
In addition, similar to LIM, PAHs can serve as tracers of the 3D density field, providing a novel probe of LSS at redshifts well beyond the reach of traditional galaxy surveys.
@@ 6.4 Future Work
The feature intensity mapping technique developed in this work is timely for the recently launched SPHEREx mission and the proposed PRIMA mission, both of which will conduct large-scale spectral-imaging surveys in the IR, where multiple PAH features fall within their wavelength coverage.
@@
We focus on PAH emission as a use case for our feature intensity mapping technique. However, this approach can be applied to mapping any broad emission feature across the electromagnetic spectrum. For example, the anomalous microwave emission has a broad spectral profile centered at GHz, which is thought to originate from spinning dust grains (C. Dickinson et al., 2018). In addition, our method can be applied to absorption spectra, such as the dust absorption feature at 2175 Å, as well as the near- and mid-infrared absorption spectra from biogenic molecular ices, including water (H2O), carbon dioxide (CO2), carbon monoxide (CO) and methanol (CH3OH) (D. Hollenbach et al., 2009; A. C. A. Boogert et al., 2011, 2013; K. I. Öberg et al., 2011).
This work is the third in our series on performing inference using the full covariance of spectral-imaging surveys, building upon Y.-T. Cheng et al. (2023) and Y.-T. Cheng et al. (2024). In Y.-T. Cheng et al. (2023), we focused on reconstructing the principal modes of the continuum from cross-correlations in broadband photometric intensity maps. In Y.-T. Cheng et al. (2024), we explored the application of the same method for line intensity mapping. In this work, we extend our model to include broad spectral features such as PAHs.
@@ 7 Conclusion
For PRIMA, we consider nine PAH features spanning to m and redshifts . We assume PHI and FIRESS surveys covering 1 and 0.1 deg2, respectively, each with 1,000 hr of integration. Our forecasts show that multiple PAH features can achieve detections of their bias-weighted intensity per bins of redshifts at , with the two bright PAH features (7.7 and 11.3 m) detectable out to .
While individually detected galaxies at high redshift will be biased toward the brightest sources, our results demonstrate that IM offers a powerful complementary approach to recover the integrated emission from the full galaxy population, including faint and unresolved sources. Our feature intensity mapping technique enables the extension of 3D IM, previously limited to narrow emission lines with LIM, to general broad spectral features such as PAHs. This is especially crucial given the importance of PAHs in tracing star formation, and dust and metallicity evolution in galaxies. The development of this technique is timely: SPHEREx has just begun its survey, and, if selected for implementation, PRIMA could launch early in the next decade.